TW6 traveling waves in opposite directions produce a standing wave. The individual wave functions are: Th6 traveling waves in opposite diretions produce a standing wave. The individual wave 4. y,-(4.0 cm) sin (3.0-2.00 y,-(4.0 cm) sin (3.0x + 2.00 where x and y are measured in centimeters. (a) Find the amplitude of the simple harmonic motion of the element of the medium located at x 2.3 cm. (b) Find the positions of the first three nodes and antinodes if one end of the string is at x ANS -4.6 em. (0, pi/3. 2pi/3) nodes, (pi/6, pi/2, 5p/6) antinodes
Answer #1

Given,

y1 = 4 cm sin(3x - 2t)

y2 = 4 cm sin(3x + 2t)

a)resulting wave will be:

y = 8 cm sn(3x) cos(wt)

for x = 2.3 ;

y = 8 cm sin(3 x 2.3) = 4.6 cm

Hence, y = 4.6 cm

b)k = 2 pi/lambda = 3 rad/cm

lambda = 2 pi/3

x = n/2 lambda = n pi.3

n = 0, 1, 2

x = 0, pi/3, 2 pi/3

And, x = (2n+1)/4 lambda

x = pi/6, pi/2,5pi/6